(x[sup]2[/sup]+3) / (x-2)(x+4)
1
Expert's answer
2011-08-25T07:58:56-0400
(x2 + 3)/(x -2)(x+4) = (x2 +3)/(x2 - 2x + 4x - 8) = (x2 +3)/(x2 + 2x - 8)